OpenStudy (anonymous):

Which of the following triangles are right triangles? Check all that apply. A. A triangle with side lengths 4, 5, 6 B. A triangle with side lengths 8, 15, 17 C. A triangle with side lengths 5, 12, 13 D. A triangle with side lengths 6 inches, 8 inches, 10 inches

OpenStudy (anonymous):

i think its d

jimthompson5910 (jim_thompson5910):

D is one answer, so you're on the right track. Are there any more answers?

jimthompson5910 (jim_thompson5910):

The key phrase "Check all that apply" means that more than one answer is possible (and most likely expected)

OpenStudy (anonymous):

umm a

jimthompson5910 (jim_thompson5910):

Let's check it... a^2 + b^2 = c^2 4^2 + 5^2 = 6^2 ... Plug in a = 4, b = 5 and c = 6 Then evaluate/simplify both sides

OpenStudy (anonymous):

Hint: Pythagoras triplets

jimthompson5910 (jim_thompson5910):

What is 4^2+5^2 ???

jimthompson5910 (jim_thompson5910):

oh wait, I just realized that you're not the OP...please let the OP respond. Thanks.

OpenStudy (anonymous):

idk arent u ssupposed to ask the one who asked the question lol

OpenStudy (anonymous):

:))

jimthompson5910 (jim_thompson5910):

yeah I'm trying that...

OpenStudy (anonymous):

op wont respod though

OpenStudy (anonymous):

i thinks shes waiting for the answer